OpenStudy (anonymous):

cos^2xcscx-cscx=___x fill in the blank

OpenStudy (anonymous):

I would bring out the common term of csc x do you know what cos^2x -1 equals? or is equivalent too?

OpenStudy (anonymous):

no....

OpenStudy (anonymous):

sin^2x + cos^2x = 1 so therefore cos^2x - 1 = sin^2x

OpenStudy (anonymous):

cscx = 1/sinx, right? so now you can simplify and have your answer - hint it will only have 1 term

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

I'm very confused

OpenStudy (anonymous):

ok..let me show you the steps

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

cscx ( cos^2x -1) = cscx (sin^2x) = 1/sinx (sin^2x) = sinx

OpenStudy (anonymous):

okay, so then my answer would be sinx?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, yeah it makes a lot more sense to see the steps, thank you :)

OpenStudy (anonymous):

it told me that it was -sin??